Calculate the pH of each of the following strong acid solutions.

4.16×10−2 M HNO3.

0.250 g of HClO3 in 1.50 L of solution.

15.00 mL of 1.10 M HCl diluted to 0.520 L .
Answer #1

7 m 1) PH = 10g (ht) HNO3 HT eta NO3- M =HA NO = 4.168102 ph=log (4.16 X10-2) = 1.381 2) moles of Heloz - map = 0.250 2.968 1

If you have any questions please comment

If you satisfied with the solution please rate it thanks